Grade 12Integral Calculus

integrate the following integral
∫ sin x sqrt(1 – cos 2x) dx

taniska
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sinx\sqrt{1-cos2x}dx
I = \int sinx\sqrt{2sin^{2}x}dx
I = \int \sqrt{2}sinx.sinxdx
I = \int \sqrt{2}sin^{2}xdx

I = \frac{1}{\sqrt{2}}\int (1-cos2x)dx
I = \frac{1}{\sqrt{2}}(x-\frac{sin2x}{2})+c